OpenStudy (anonymous):

Prove by induction 2n + 3 ≤ 2^n for all integers n ≥ N. You should find the minimal integer value of N for which this proposition is true and then construct the proof using this value.

OpenStudy (anonymous):

please?

OpenStudy (anonymous):

probably true if \(n=4\) right ?

OpenStudy (anonymous):

well actually since it is \(\leq \) it is true if \(n=3\)

OpenStudy (anonymous):

N = 4 not 3

OpenStudy (anonymous):

satellite73, can't believe you are still here, you are the one who helped me pass analysis haha! i haven't done induction in a while so i forgot how to do this especially since they are teaching us to only use one side of the equation in this module. Yes N would be 4

OpenStudy (anonymous):

i am glad you passed

OpenStudy (anonymous):

assume true for \(k\) that is assume \(2k + 3 ≤ 2^k\) now lets try it for \(n=k+1\) you get \[2(k+1)+3=2k+5\] and by induction \(2k+3<2^k\) so \(2k+5<2^k+2<2^{k+1}\)

OpenStudy (anonymous):

i don't understand the last line

OpenStudy (anonymous):

The so bit

OpenStudy (anonymous):

Now prove 2^k + 2 <= 2^(k+1)... note n >= 4 so 2^n >= 16

OpenStudy (anonymous):

where is the 2^k + 2 coming from

OpenStudy (anonymous):

2k + 3 < 2^k therefore 2k + 5 < 2^k + 2; the goal is to prove 2k + 5 < 2^(k+1)

OpenStudy (anonymous):

<= ***

OpenStudy (anonymous):

so if we know that 2k + 5 < 2^k + 2 is true and we have to prove that 2k + 5 < 2^(k+1) is true then do i have to prove that 2^k+2=2^(k+1) am i understanding this correctly? haven't done induction with inequalities before this is a total mess :/

OpenStudy (anonymous):

No, you want to prove 2^k + 2 <= 2^(k+1) :-) If k >= 4, then 2^k >= 16, right? Therefore 2^k + 2 <= 18 <= 32 <= 2^(k+1)

OpenStudy (anonymous):

thanks for your help